OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 8(1.05)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.26 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 2 to n = 6, and what does it represent? (4 points)

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@PeterPan

OpenStudy (anonymous):

\[f(n) = 8(1.05)^n \] When the scientist concluded his study, the height of the plant was approximately 11.26 cm. What is a reasonable domain to plot the growth function? \[11.26=8\times (1.05)^n\] solve for \(n\)

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

i dunno i didn't do it hold on and i can check

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

divide by 8, use the change of base formula and a calculator of course since you have to use a calculator anyways you might as well cut to the chase http://www.wolframalpha.com/input/?i=11.26%3D8%281.05%29^x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it would be 7?

OpenStudy (anonymous):

that is what the wolf says, yes wolf is very smart so i would not argue

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

that would put the domain at \(0\leq n\leq 7\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What does the y-intercept of the graph of the function f(n) represent? what is what you get if \(n=0\) which is how high it was at the beginning, when he started counting

OpenStudy (anonymous):

it is the number out front since \(b^0=1\) always

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What is the average rate of change of the function f(n) from n = 2 to n = 6, \[\frac{f(6)-f(2)}{6-2}\] whatever that is

OpenStudy (anonymous):

and what does it represent? (4 points) it represents... the average rate of change! (doh)

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Can you help with part b @satellite73